# SOMONITOR: Explainable Marketing Data Processing and Analysis with Large Language Models

### v1

@@ 7. Conclusion
In this work, we have introduced an innovative explainable AI framework SOMONITOR that aims to combine human intuition with the efficiency of modern AI to significantly enhance the capabilities of marketers across all stages of the marketing funnel, from strategic planning to content creation and campaign execution. The challenges traditionally faced by marketers, especially the need to manually sift through immense quantities of transient online content, are addressed by SOMONITOR’s ability to automate and refine the process of competitor analysis, content research, and strategic branding. SOMONITOR ranks advertising content by its potential to achieve higher click-through rates via a specially trained model and then employs LLMs to distill core content pillars such as target audiences, customer needs, and product features from successful competitor marketing endeavors. These pillars then lead to communication themes and targeted customer personas, providing a structured basis for content and campaign planning. By integrating these insights with data from a brand’s own marketing campaigns, SOMONITOR not only constructs compelling narratives for new customer personas but also generates detailed content briefs in the form of customer persona stories, empowering marketing teams to produce content and execute campaigns. We believe that SOMONITOR can transform digital marketing operations, enabling marketers to quickly navigate through large datasets and derive actionable insights that boost campaign effectiveness and enhance job satisfaction, redefining traditional marketing approaches (Farseev, 2024).

### v2

@@ X Conclusion
In this work, we have introduced an innovative explainable AI framework SOMONITOR that aims to combine human intuition with the efficiency of modern AI to significantly enhance the capabilities of marketers across all stages of the marketing funnel, from strategic planning to content creation and campaign execution. The challenges traditionally faced by marketers, especially the need to manually sift through immense quantities of transient online content, are addressed by SOMONITOR’s ability to automate and refine the process of competitor analysis, content research, and strategic branding, which has been proven by a case study in the real-world settings. SOMONITOR ranks advertising content by its potential to achieve higher click-through rates via a specially trained model and then employs LLMs to distill core content pillars, which lead to communication themes and targeted customer personas, providing a structured basis for content and campaign planning. By integrating these insights with brand marketing data, SOMONITOR not only constructs compelling narratives for new customer personas but also generates detailed content briefs in the form of customer persona stories, empowering marketing teams to produce content and execute campaigns in a guided manner, which has been proven to help in prioritizing different content elements and enhance real-world advertising performnace. We believe that SOMONITOR can transform digital marketing operations, enabling marketers to quickly navigate through large datasets and derive actionable insights that boost campaign effectiveness and enhance job satisfaction, redefining traditional marketing approaches [111].
XI Acknowledgement
This work was funded by the Russian Science Foundation grant №. 22-11-00135, https://rscf.ru/en/project/22-11-00135/